OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

3y=x+8

OpenStudy (lgbasallote):

use this equation \[\huge y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\]